a dealer wishes to produce a 60 pound mixture of candy to sell at $5 a pound< by mixing candy of $7 a pound and candy of $4 a pound. how many pound of each should u use?
Answer by lwsshak3(11628)

a dealer wishes to produce a 60 pound mixture of candy to sell at $5 a pound< by mixing candy of $7 a pound and candy of $4 a pound. how many pound of each should u use?
**
let x= pounds of $7 canday used for mixture
60- x= pounds of $4 canday used for mixture
..
7x+4(60-x)=5(60)
7x+240-4x=300
3x=60
x=20
60-20=40 pounds
ans:
You should use:
20 lbs of $7 candy and
40 lbs of $4 candy